how would i graph y=x+2

Answer by richwmiller(17219)

How would I graph y=x+2
Since a line can be made from two points all you need to do is
choose two (or more) x's and y's and plot them
The easiest two points to plot and calculate are when x=0 and when y=0
if x=0 then y=2
if y=0 then x=-2(minus 2)
Or you can just pick any x and y of your choice and plot them
for example
if x=1 then y=1+2=3
if x=4 then y=4+2=6